# bibdata.bib

### v1

@@ V-B2 Experiments on EPFL Benchmarks
V-C Experiments Under the MaxHD Constraint
This set of experiments approximates the ISCAS85 benchmarks in Table I under the MaxHD constraint. We compare our ALS flow with MECALS. We modify the open-source codes of MECALS to support the MaxHD constraint and run both methods on the same platform for fair comparison. MUSCAT is not compared in this experiment as its open-source code does not support the MaxHD constraint. The MaxHD bounds are set to $\lfloor 0.1O\rfloor$ and $\lfloor 0.2O\rfloor$ for each benchmark, where $O$ is the number of POs of the benchmark. The same types of LACs, i.e., the constant LAC [shin2011new] and the SASIMI LAC [venkataramani2013substitute], are used in our method and MECALS in this experiment.
Table IV compares the area ratio, delay ratio, and runtime of our flow and MECALS on the ISCAS85 benchmarks under the MaxHD constraint. We can see that our method achieves smaller area and delay ratios on most benchmarks. Our flow can achieve an average area ratio of $63.0\%$ and an average delay ratio of $73.5\%$ with an average runtime of $12$ seconds. Compared to MECALS, our method reduces area and delay by $12.6\%$ and $10.5\%$, respectively, on average. Besides, our method reduces area and delay in comparison to MECALS for most benchmarks, and there is a significant improvement in area saving of more than $50\%$ for the benchmark c7552. Although for the benchmark c1908 under the MaxHD bound of $5$ and the benchmark c2670 under the MaxHD bound of $28$, our method is worse than MECALS in terms of the area ratio, but our method reduces more delay. Moreover, our method is more efficient than MECALS in all benchmarks except for the benchmark c2670 under the MaxHD bound of $14$, accelerating over MECALS by 13.6×13.6$\times$13.6 × on average.
V-D Effectiveness of Simulation-Guided LAC Pruning
To show the effectiveness of the simulation-guided LAC pruning (see Section IV-B) in our ALS flow, we conduct an ablation study on the arithmetic circuits used in MECALS in Table I under the MaxED constraint. We compare our flows with and without the simulation-guided LAC pruning. For the flow without the pruning, we evaluate all candidate LACs in each iteration and select the first $K$ valid LACs to simplify the circuit. Similar to the previous experiments, we choose MaxED bounds of $\lfloor 2^{0.1O}\rfloor$ and $\lfloor 2^{0.2O}\rfloor$ for each benchmark, where $O$ is the number of POs of the benchmark. Since the flow without the pruning is very slow, we only use the simple constant LACs in this experiment to ensure that the experiment finishes in a reasonable time.
@@ V-D Effectiveness of Simulation-Guided LAC Pruning
Figure 5: Comparison between the approximate adders and multipliers synthesized by our flow and those from the EvoApproxLib under the MaxED constraint.
@@ V-E Comparison on Approximate Adder and Multiplier Designs
Given the importance of approximate adders and multipliers, especially in AI hardware accelerators, we compare the approximate adders and multipliers synthesized by our ALS flow with those from the EvoApproxLib [mrazek2017evoapprox8b, mrazek2022evo] (version 2022), a widely-used open-source library of approximate adders and multipliers. The benchmarks compared are the 12-bit and 16-bit unsigned adders and the 8-bit, 11-bit, 12-bit, and 16-bit unsigned multipliers. The approximate designs from the EvoApproxLib are those synthesized under the MaxED constraint. Our ALS flow starts from the accurate circuits from the EvoApproxLib, converts them to AIGs, and applies the constant and SASIMI LACs to simplify the circuits under the MaxED constraint. The MaxED bounds are set as the MaxEDs of the approximate circuits from the EvoApproxLib, which means that the approximate circuits generated by our flow have the same or smaller MaxEDs compared to those in the EvoApproxLib.
@@ Acknowledgement
This work is supported by the Swiss National Science Foundation Grant “Supercool: Design methods and tools for superconducting electronics” with funding number 200021_1920981, by the Synopsys Inc., and by the National Natural Science Foundation of China under Grants T2293700, T2293701, and T2293704.
[…]
[Uncaptioned image] Chang Meng is a postdoctoral researcher at the Integrated Systems Laboratory, EPFL Lausanne, Switzerland. He received his Ph.D. degree in Electronic Science and Technology at Shanghai Jiao Tong University in 2023. His research interest is electronic design automation for emerging computing paradigms, especially the logic synthesis and verification of approximate computing circuits. His research work was nominated for the Best Paper Award at Design, Automation, and Test in Europe Conference (DATE).
[Uncaptioned image] Weikang Qian is an associate professor in the University of Michigan-Shanghai Jiao Tong University Joint Institute at Shanghai Jiao Tong University. He received his Ph.D. degree in Electrical Engineering at the University of Minnesota in 2011 and his B.Eng. degree in Automation at Tsinghua University in 2006. His main research interests include electronic design automation and digital design for emerging computing paradigms. His research works were nominated for the Best Paper Awards at International Conference on Computer-Aided Design (ICCAD), Design, Automation, and Test in Europe Conference (DATE), and International Workshop on Logic and Synthesis (IWLS). He serves as an associate editor of the IEEE Transactions on Computer-Aided Design of Integrated Circuits and Systems. He is a senior member of IEEE.
[Uncaptioned image] Giovanni De Micheli is Professor and Director of the Integrated Systems Laboratory and Scientific Director of the EcoCloud center at EPFL Lausanne, Switzerland. Previously, he was Professor of Electrical Engineering at Stanford University. He holds a Nuclear Engineer degree (Politecnico di Milano, 1979), a M.S. and a Ph.D. degree in Electrical Engineering and Computer Science (University of California at Berkeley, 1980 and 1983). He is a Fellow of ACM, AAAS and IEEE, a member of the Academia Europaea, of the Swiss Academy of Engineering Sciences, and International Honorary member of the American Academy of Arts and Sciences. His current research interests include several aspects of design technologies for integrated circuits and systems, such as synthesis for emerging technologies. He is also interested in heterogeneous platform design including electrical components and biosensors, as well as in data processing of biomedical information. He is member of the Scientific Advisory Board of IMEC (Leuven, B) and STMicroelectronics. Professor De Micheli is the recipient of the 2025 IEEE Gustav Kirchhoff Award, the 2022 ESDA-IEEE/CEDA Phil Kaufman Award, the 2019 ACM/SIGDA Pioneering Achievement Award, and several other awards.

### v2

@@ V-D Effectiveness of Simulation and Parameter Study
V-D1 Effectiveness of Simulation-Guided LAC Pruning
To show the effectiveness of the simulation-guided LAC pruning (see Section IV-B) in our ALS flow, we conduct an ablation study on the arithmetic circuits used in MECALS in Table I under the MaxED constraint. We compare our flows with and without the simulation-guided LAC pruning. For the flow without the pruning, we evaluate all candidate LACs in each iteration and select the first $K$ valid LACs to simplify the circuit. Similar to the previous experiments, we choose MaxED bounds of $\lfloor 2^{O/10}\rfloor$ and $\lfloor 2^{O/5}\rfloor$ for each benchmark, where $O$ is the number of POs of the benchmark. Since the flow without the pruning is very slow, we only use the simple constant LACs in this experiment to ensure that the experiment finishes in a reasonable time.
@@ V-D1 Effectiveness of Simulation-Guided LAC Pruning
Figure 5: Impact of simulation count on final area after applying our method and runtime of our method under $\lfloor 2^{O/5}\rfloor$ MaxED constraint. Runtime data differs from that in Table II since a different computer is used in this experiment.
V-D2 Impact of Simulation Count on Synthesis Quality and Runtime
We study the impact of the simulation count ($M_{\textit{small}}$ and $M$ in Section IV-B1) on both synthesis quality and runtime. Recall that our method adopts a two-phase simulation-guided LAC pruning (see Section IV-B1). First, $M_{\textit{small}}$ patterns are used for rough pruning of LACs, followed by a larger set of $M$ patterns for further filtering. We evaluate the impact of $M$ and $M_{\textit{small}}$ on four benchmarks, binsqrd, buttfly, mult16, and mult32, under a fixed MaxED constraint of $\lfloor 2^{O/5}\rfloor$, where $O$ is the PO number of the benchmark. For each benchmark, we test seven $(M_{\textit{small}},M)$ pairs: $(2^{5},2^{8})$, $(2^{5},2^{11})$, $(2^{5},2^{13})$, $(2^{5},2^{16})$, $(2^{10},2^{11})$, $(2^{10},2^{13})$, and $(2^{10},2^{16})$. Note that $(2^{10},2^{13})$ is the default setting in all previous experiments.
Fig. 5 shows the impact of the simulation count on the final circuit area and runtime of our method. For all benchmarks, under the same $M_{\textit{small}}$, the final area decreases or remains unchanged as $M$ increases. This is because a larger $M$ can filter out more invalid LACs, thereby retaining more valid ones in the top $K$ candidates for further evaluation, which produces smaller approximate circuits. Meanwhile, under the same $M_{\textit{small}}$, the runtime generally increases with $M$, since more simulation patterns require more simulation time. An exception occurs for mult32 with $M_{\textit{small}}=2^{5}$, where runtime first decreases and then increases. When both $M_{\textit{small}}$ and $M$ are small, pruning is ineffective, and many invalid LACs remain in the top $K$ LACs and must be verified by SAT solving, which dominates the runtime. As $M$ increases from $2^{8}$ to $2^{13}$, more invalid LACs are eliminated by simulation, leading to fewer SAT problems and reduced runtime. When $M$ further increases to $2^{16}$, the runtime rises again due to the long simulation time. Overall, the default setting of $(M_{\textit{small}},M)=(2^{10},2^{13})$, shown as brown bars in Fig. 5, provides a good trade-off between the final area and runtime across the four benchmarks.
V-D3 Impact of Simulation Count on SAT Solving Statistics
We further study how the simulation count influences the SAT solving statistics. In this study, we fix $M_{\textit{small}}=2^{5}$ and vary $M\in\{2^{8},2^{11},2^{13},2^{16}\}$, while keeping the other settings identical to those in the previous experiment in Section V-D2. Table VI shows, for different $M$ values, the total number of SAT problems solved, the percentages of SAT and UNSAT results, and the number of UNDEFINED results observed in the first two iterations of our method. Since different $M$ values lead to different iteration numbers, we restrict the comparison to the first two iterations for fairness. As $M$ increases, the percentage of SAT results generally decreases while that of UNSAT results generally increases. This trend is expected because a larger $M$ prunes more invalid LACs, which increases the likelihood that the preserved top $K$ LACs are valid. As valid LACs correspond to UNSAT results when their maximum errors are formally checked, the UNSAT ratio rises, while the SAT ratio decreases. Moreover, UNDEFINED results are only observed in mult32 when $M=2^{11},2^{13},$ and $2^{16}$. Due to the large size of mult32, the SAT solver may fail to resolve some instances within the computation budget.
TABLE VI: Impact of simulation count on SAT solving statistics in the first two iterations of our method under $\lfloor 2^{O/5}\rfloor$ MaxED constraint. $M_{\textit{small}}$ is fixed to $2^{5}$.
Circuit $M$ All SAT calls SAT percentage UNSAT percentage #UNDEF count
binsqrd $2^{8}$ 98 42.9% 57.1% 0
$2^{11}$ 60 3.3% 96.7% 0
$2^{13}$ 60 3.3% 96.7% 0
$2^{16}$ 60 3.3% 96.7% 0
buttfly $2^{8}$ 100 95.0% 5.0% 0
$2^{11}$ 100 95.0% 5.0% 0
$2^{13}$ 100 95.0% 5.0% 0
$2^{16}$ 36 80.6% 19.4% 0
mult16 $2^{8}$ 17 17.6% 82.4% 0
$2^{11}$ 15 0.0% 100.0% 0
$2^{13}$ 15 0.0% 100.0% 0
$2^{16}$ 15 0.0% 100.0% 0
mult32 $2^{8}$ 106 42.5% 57.5% 0
$2^{11}$ 69 5.8% 88.4% 4
$2^{13}$ 33 6.1% 90.9% 1
$2^{16}$ 33 6.1% 90.9% 1
TABLE VII: Impact of number of top LACs selected ($K$) on final area, runtime, and number of iterations of our method under $\lfloor 2^{O/5}\rfloor$ MaxED constraint. Runtime data differs from that in Table II since a different computer is used in this experiment.
[…]
binsqrd 50 801.7 562 10
100 797.5 445 7
1000 795.6 851 5
10000 790.3 927 5
buttfly 50 127.2 48 7
100 127.2 73 7
1000 127.2 49 5
10000 127.2 32 4
mult16 50 1327.3 149 12
100 1316.2 142 7
1000 1323.1 123 5
10000 1328.9 729 4
mult32 50 5379.1 1240 19
100 5353.0 866 10
1000 5360.4 1377 7
10000 5382.2 2924 4
V-D4 Impact of Number of Top LACs (Parameter $K$)
Recall that during the LAC selection step shown in Fig. 4, only the top $K$ LACs are kept for further evaluation based on their maximum error lower bounds and estimated area reductions. We study the impact of $K$ on the final area, runtime, and iteration number of our method. Specifically, we test $K\in\{50,100,1000,10000\}$ on the four benchmarks binsqrd, buttfly, mult16, and mult32 under the MaxED constraint of $\lfloor 2^{O/5}\rfloor$. The results are shown in Table VII.
The impact of $K$ on the final area and runtime varies across benchmarks, but a common related trend can be observed. As $K$ increases, the number of iterations decreases, since more candidate LACs are evaluated and applied in each iteration. However, a larger $K$ with fewer iterations does not necessarily yield shorter runtime, because evaluating more LACs per iteration also requires solving more SAT problems per iteration. Thus, the overall runtime is determined by the trade-off between the iteration count and the number of SAT calls per iteration. For binsqrd, mult16, and mult32, runtime decreases at first as $K$ increases, but then rises again when $K$ further increases. For buttfly, runtime exhibits the opposite trend, first increasing and then decreasing.
The effect of $K$ on the final circuit area also differs by benchmarks. For binsqrd, the final area consistently decreases with $K$. For buttfly, the final area remains the same for different $K$ values. For mult16 and mult32, the final area first decreases and then increases as $K$ grows. One possible reason is that a small $K$ may exclude promising LACs from further evaluation, leading to suboptimal results with larger area. Conversely, when $K$ is very large, too many LACs are applied in each iteration under the greedy selection strategy (see Fig. 4). This causes the circuit to deviate significantly from the current approximate circuit, making the previously computed maximum error lower bounds of LACs inaccurate. Guided by these inaccurate values, some poor LACs with large errors may be applied, which eventually causes the flow to terminate prematurely and produces a circuit with larger area.
@@ V-D4 Impact of Number of Top LACs (Parameter $K$)
Figure 6: Comparison between approximate adders and multipliers synthesized by our flow and those from the EvoApproxLib under the MaxED constraint.
@@ V-E Comparison on Approximate Adder and Multiplier Designs
Given the importance of approximate adders and multipliers, especially in AI accelerators, we compare the approximate adders and multipliers synthesized by our ALS flow with those from the EvoApproxLib [22] (version 2022), a widely-used open-source library of approximate adders and multipliers. The benchmarks compared are the 12-bit and 16-bit unsigned adders and the 8-bit, 11-bit, 12-bit, and 16-bit unsigned multipliers. The approximate designs from the EvoApproxLib are those synthesized under the MaxED constraint. Our ALS flow starts from the accurate circuits from the EvoApproxLib, converts them to AIGs, and applies the constant and SASIMI LACs to simplify the circuits under the MaxED constraint. The MaxED bounds are set as the MaxEDs of the approximate circuits from the EvoApproxLib, which means that the approximate circuits generated by our flow have the same or smaller MaxEDs compared to those in the EvoApproxLib.
@@ Acknowledgement
This work is supported in part by the Synopsys Inc., by the National Natural Science Foundation of China under Grant 62574132, and by the Natural Science Foundation of Shanghai under Grant 25ZR1401189. We thank the anonymous reviewers for their constructive feedback.